# Do Not Design, Learn: A Trainable Scoring Function for Uncertainty Estimation in Generative LLMs

### v1

@@ 5.1 Experimental Setup
To test the performance of UE methods, we employ 3 different closed-ended QA datasets. Following (Kuhn et al., 2023), we use a subset of the validation set of TriviaQA Joshi et al. (2017). Second, we test on the entire validation split of NaturalQA Kwiatkowski et al. (2019). Lastly, we combine train and validation splits of Web Questions, shortly WebQA Berant et al. (2013).
Dataset UE Method Scoring Function Llama2-7b Llama3-8b Mistral-7b Gemma-7b
TriviaQA Lexical Similarity - 0.647 0.683 0.720 0.594
# Semantic Groups - 0.792 0.819 0.757 0.728
$\mathbf{p}$(True) - 0.616 0.842 0.808 0.713
Confidence LNS 0.697 0.748 0.722 0.604
MARS 0.751 0.799 0.745 0.602
TokenSAR 0.747 0.792 0.747 0.604
LARS 0.851 0.872 0.844 0.835
Entropy LNS 0.692 0.747 0.738 0.596
MARS 0.736 0.801 0.755 0.600
TokenSAR 0.734 0.793 0.763 0.605
LARS 0.842 0.864 0.849 0.830
SE LNS 0.795 0.835 0.810 0.732
MARS 0.797 0.845 0.810 0.729
TokenSAR 0.796 0.839 0.813 0.729
LARS 0.849 0.866 0.854 0.828
NaturalQA Lexical Similarity - 0.600 0.651 0.637 0.546
# Semantic Groups - 0.705 0.736 0.675 0.656
$\mathbf{p}$(True) - 0.561 0.761 0.730 0.683
Confidence LNS 0.677 0.697 0.666 0.608
MARS 0.714 0.717 0.692 0.645
TokenSAR 0.703 0.717 0.682 0.637
LARS 0.780 0.812 0.782 0.784
Entropy LNS 0.661 0.698 0.679 0.597
MARS 0.707 0.707 0.701 0.646
TokenSAR 0.683 0.714 0.694 0.633
LARS 0.775 0.805 0.781 0.779
SE LNS 0.721 0.759 0.727 0.667
MARS 0.730 0.750 0.735 0.670
TokenSAR 0.721 0.756 0.726 0.669
LARS 0.772 0.794 0.778 0.785
Table 1: AUROC performance of UE methods.
@@
To train the model of the proposed method LARS, we employ subsets of the train splits of TriviaQA Joshi et al. (2017) and NaturalQA Kwiatkowski et al. (2019). We randomly select $\sim$13k questions from each dataset and sample six generations per question, ensuring the most likely generation is included, for each model mentioned below. From these generations, we curate unique QA pairs for calibration data. Typically, we train distinct LARS models for each model-dataset combination. In some experiments, we merge TriviaQA and NaturalQA for each model and train accordingly, which we specify when used. To obtain binary ground truths for QA pairs, we utilize GPT-3.5-turbo as in Bakman et al. (2024); Lin et al. (2023); Chen and Mueller (2023). Please refer to Appendix D for details and prompt.
Models.
We test UE methods on 4 popular models. Llama2-7b-chat Touvron et al. (2023) and Llama3-8b-instruct AI@Meta (2024) are optimized for dialogue use cases. Mistral-7b-instruct Jiang et al. (2023) and Gemma-7b-it Team et al. (2024) are instruction tuned versions of the corresponding base models. For the sake of simplicity, we do not use instruction indicator words of the models in the rest of the paper.
Metrics.
Following previous works, we calculate AUROC (Area Under the Receiver Operating Characteristic) score, a commonly used metric used to evaluate the performance of a binary classifier Kuhn et al. (2023); Bakman et al. (2024); Duan et al. (2024). The ROC curve plots the true positive rate against the false positive rate at various thresholds. AUROC score is the area under this curve, and it provides a single number that summarizes the model’s ability to discriminate between the positive and negative classes regardless of the threshold. An AUROC score of 1.0 represents a perfect classifier, while 0.5 is equivalent to random guessing.
Baselines.
We use three probability-based UE methods following (Bakman et al., 2024). Confidence is the negative of the response score. It is calculated as the negative score of the most likely generation to a given question. The other UE methods are Entropy as in (2) and Semantic Entropy (SE) (3). Each method uses a scoring function to assign a score to a model generation. We compare LARS with 3 SOTA scoring functions for this purpose: Length-normalized scoring (LNS)Malinin and Gales (2021), MARS Bakman et al. (2024) and TokenSAR Duan et al. (2024). Our proposal LARS is a scoring function, compared with other baseline scoring functions combined with all probability-based UE methods.
Further, We add three non-probability-based UE approaches to our baseline set. Lexical Similarity Fomicheva et al. (2020), is the average of the Rouge-L scores between unique sampled generation pairs to a given question. $p$(True) Kadavath et al. (2022), a self-check method, asks the model itself if the most likely answer is correct by providing the question, sampled generations, and the answer. Lastly, following Kuhn et al. (2023), we compare with # Semantic Groups, the number of semantic clusters, as in SE. In all of our experiments, number of sampled generations is 5.
UE Method Scoring Function Llama2-7b Llama3-8b Mistral-7b Gemma-7b
Lexical Similarity - 0.643 0.640 0.645 0.607
# Semantic Groups - 0.612 0.599 0.601 0.594
$\mathbf{p}$(True) - 0.558 0.636 0.668 0.677
Confidence LNS 0.656 0.645 0.634 0.608
MARS 0.669 0.659 0.637 0.607
TokenSAR 0.664 0.656 0.640 0.607
LARS (TriviaQA only) 0.718 0.704 0.681 0.739
LARS (NaturalQA only) 0.701 0.690 0.682 0.756
LARS (TriviaQA+NaturalQA) 0.715 0.713 0.686 0.739
Entropy LNS 0.656 0.650 0.647 0.610
MARS 0.675 0.664 0.647 0.616
TokenSAR 0.668 0.661 0.649 0.610
LARS (TriviaQA only) 0.719 0.704 0.690 0.730
LARS (NaturalQA only) 0.712 0.690 0.691 0.748
LARS (TriviaQA+NaturalQA) 0.714 0.703 0.693 0.733
SE LNS 0.672 0.664 0.665 0.629
MARS 0.679 0.669 0.665 0.629
TokenSAR 0.674 0.667 0.663 0.625
LARS (TriviaQA only) 0.716 0.697 0.689 0.732
LARS (NaturalQA only) 0.709 0.685 0.693 0.745
LARS (TriviaQA+NaturalQA) 0.711 0.694 0.697 0.729
Table 2: AUROC performance of UE methods with different scoring functions on WebQA dataset. LARS models are trained with TriviaQA and/or NaturalQA.
@@ 5.2 Main Results
We present the results of our method alongside other baselines in Table 1. Notably, LARS significantly enhances the performance of all existing scoring functions across each probability-based UE method, with improvements reaching up to 0.231 points over LNS. Additionally, LARS boosts the confidence metric to levels comparable with Semantic Entropy (SE) and Entropy. This is particularly important considering the inference cost: Entropy-based methods require multiple output samples (5 in our experiments), which can be computationally expensive in the context of LLMs. Further, SE necessitates $O(N^{2})$ model passes for semantic clustering, where $N$ is the number of sampled outputs. In contrast, LARS operates with a single pass using a RoBERTa-based model with 125M parameters—a computation level that is negligible compared to models with capacities of 7B parameters or more. Lastly, the LARS scoring function demonstrates that probability-based UE methods outperform response clustering methods, including Lexical Similarity, the number of Semantic Groups, and the self-checking method $p(\text{True})$.
@@ 5.3 Out-of-Distribution (OOD) Experiments
First, we investigate how the performance of LARS is affected when the model encounters questions which have a distribution deviating from that of the calibration set. To this end, we conduct tests using WebQA, with LARS models trained on TriviaQA and/or NaturalQA for each distinct chat model. The results are presented in Table 2, and additional results on out-of-distribution (OOD) data generalization are available in Appendix C.2. Impressively, LARS, despite being trained on different datasets, outperforms all other scoring functions across all probability-based UE methods, achieving an average improvement of approximately $\sim 0.04$ points.
@@
Next, we analyze how LARS performs when the responses in the calibration set are derived from a different chat model than the one used at test time. Due to space limitations, we provide a subset of the results in Table 3; however, comprehensive results are presented in Appendix C.1. Notably, optimal LARS performance is achieved when the same chat model is used for both training and testing. Nevertheless, OOD model scores still surpass those of baseline scoring functions (see Table 1 for baselines), confirming the effectiveness of LARS.
@@
Confidence Llama2-7b 0.858 0.852 0.835
Llama3-8b 0.836 0.874 0.833
Mistral-7b 0.831 0.850 0.852
Entropy Llama2-7b 0.842 0.852 0.841
Llama3-8b 0.823 0.864 0.841
Mistral-7b 0.827 0.850 0.849
SE Llama2-7b 0.850 0.863 0.850
Llama3-8b 0.836 0.872 0.849
Mistral-7b 0.840 0.862 0.859
@@
Figure 3: AUROC scores of two different probability association methods for LARS on 2 datasets and 4 models.
5.4 Turkish TriviQA Experiment
To experimentally support our claims regarding the limitations of existing scoring functions in low-resource languages discussed in Section 3, we translated the TriviaQA test and calibration datasets into Turkish using the Googletrans  [^1] https://py-googletrans.readthedocs.io. As illustrated in Table 4, the performance gains of MARS and TokenSAR over the LNS baseline are diminished in the Turkish dataset. This decline is particularly notable for MARS, which incorporates language-specific assumptions in its design, such as phrase separation. In contrast, LARS continues to demonstrate a significant advantage, maintaining its superiority even though the RoBERTa model is pretrained in English. This indicates that calibration training enables LARS to adapt effectively to different languages.
Scoring Function English Turkish
LNS 0.747 0.692
MARS 0.791 (+0.044) 0.695 (+0.003)
TokenSAR 0.793 (+0.046) 0.720 (+0.028)
LARS 0.864 (+0.117) 0.814 (+0.122)

### v2

@@ 5.1 Experimental Setup
To evaluate UE methods, we use a mathematical reasoning dataset and three closed-ended QA datasets. Specifically, we utilize the complete test set of GSM8K for mathematical reasoning Cobbe et al. (2021). Following (Kuhn et al., 2023), we select a subset of the validation set from TriviaQA Joshi et al. (2017). Additionally, we evaluate using the entire validation split of NaturalQA Kwiatkowski et al. (2019). Finally, we combine the training and validation splits of Web Questions (WebQA) Berant et al. (2013).
Models.
We test UE methods on 5 popular open-weight models. Llama2-7b-chat, Llama2-13b-chat Touvron et al. (2023) and Llama3-8b-instruct AI@Meta (2024) are optimized for dialogue use cases. Mistral-7b-instruct Jiang et al. (2023) and Gemma-7b-it Mesnard et al. (2024) are instruction tuned versions of the corresponding base models. For the sake of simplicity, we drop instruction indicator words such as "-chat" in the rest of the paper.
Metrics.
Following previous works, we set the model’s golden (most-probable) generation’s correctness [^1] With given ground truth and model generation, we use GPT-3.5-turbo for evaluating the correctness of the generation Lin et al. (2024); Duan et al. (2024); Bakman et al. (2024) as labels (0 and 1) and UE scores as predictions Kuhn et al. (2023); Bakman et al. (2024); Duan et al. (2024). Using this, we calculate the AUROC (Area Under the Receiver Operating Characteristic), a common metric for binary classifiers Kuhn et al. (2023); Duan et al. (2024); Lin et al. (2024). Since AUROC is sensitive to data imbalance, we also include the Prediction Rejection Ratio (PRR) Malinin and Gales (2021). AUROC scores range from 0.5 (random) to 1.0 (perfect), and PRR ranges from 0.0 (random) to 1.0 (perfect).
Baselines.
We use 4 probability-based UE methods. Confidence is calculated as the negative of the most likely generation’s score for a given question. The other UE methods are Entropy as in (2), Semantic Entropy (SE) as in (3), and SentSAR Duan et al. (2024). Each method employs a scoring function to assign a score to a generation. We compare LARS with 3 SOTA scoring functions for this purpose: Length-Normalized Scoring (LNS) Malinin and Gales (2021), MARS Bakman et al. (2024) and TokenSAR Duan et al. (2024). LARS is evaluated against these scoring functions across all probability-based UE methods. It is worth noting that combining SentSAR and TokenSAR results in the SAR method Duan et al. (2024).
UE Method Scoring Llama2-7b Llama3-8b Mistral-7b Gemma-7b Llama2-13b
Function AUROC PRR AUROC PRR AUROC PRR AUROC PRR AUROC PRR
WebQA Lex. Sim. - 0.643 0.310 0.640 0.321 0.645 0.312 0.608 0.214 0.624 0.261
# Sem. Gr. - 0.612 0.138 0.599 0.143 0.601 0.184 0.630 0.213 0.587 0.157
$\mathbf{p}$(True) - 0.558 0.078 0.636 0.290 0.667 0.358 0.552 0.041 0.580 0.171
Eccentricity - 0.680 0.375 0.674 0.386 0.662 0.333 0.606 0.203 0.686 0.358
Degree Matrix - 0.683 0.380 0.676 0.384 0.662 0.326 0.611 0.195 0.682 0.364
Confidence LNS 0.656 0.329 0.645 0.324 0.634 0.305 0.625 0.246 0.602 0.233
MARS 0.669 0.349 0.649 0.333 0.637 0.316 0.627 0.258 0.585 0.199
TokenSAR 0.664 0.345 0.656 0.347 0.640 0.320 0.657 0.287 0.615 0.248
LARS (OOD) 0.715 0.430 0.713 0.464 0.686 0.406 0.726 0.442 0.676 0.367
SE LNS 0.672 0.360 0.664 0.366 0.665 0.353 0.675 0.334 0.644 0.297
MARS 0.679 0.367 0.667 0.370 0.665 0.354 0.679 0.340 0.632 0.267
TokenSAR 0.674 0.365 0.667 0.372 0.663 0.351 0.680 0.343 0.647 0.298
LARS (OOD) 0.711 0.440 0.694 0.449 0.697 0.430 0.719 0.440 0.678 0.382
GSM8K Lex. Sim. - 0.444 0.000 0.632 0.272 0.537 0.019 0.544 0.080 0.551 0.110
# Sem. Gr. - 0.513 0.000 0.584 0.138 0.532 0.037 0.566 0.114 0.561 0.065
$\mathbf{p}$(True) - 0.540 0.099 0.797 0.623 0.665 0.238 0.486 0.000 0.501 0.000
Eccentricity - 0.547 0.049 0.664 0.384 0.584 0.109 0.595 0.146 0.600 0.163
Degree Matrix - 0.535 0.056 0.667 0.667 0.604 0.165 0.584 0.117 0.605 0.179
Confidence LNS 0.570 0.031 0.686 0.390 0.567 0.072 0.556 0.370 0.615 0.196
MARS 0.567 0.010 0.713 0.438 0.568 0.076 0.541 0.099 0.562 0.114
TokenSAR 0.579 0.045 0.719 0.460 0.619 0.156 0.579 0.161 0.636 0.233
LARS 0.720 0.319 0.836 0.711 0.708 0.350 0.706 0.370 0.738 0.497
LARS (OOD) 0.603 0.097 0.684 0.348 0.630 0.188 0.576 0.114 0.635 0.218
SE LNS 0.516 0.000 0.633 0.321 0.560 0.076 0.588 0.141 0.587 0.153
MARS 0.513 0.000 0.640 0.344 0.563 0.080 0.586 0.134 0.583 0.122
TokenSAR 0.526 0.005 0.638 0.344 0.578 0.102 0.588 0.148 0.592 0.171
LARS 0.675 0.267 0.715 0.528 0.663 0.310 0.679 0.345 0.697 0.383
LARS (OOD) 0.572 0.072 0.633 0.298 0.605 0.170 0.579 0.112 0.608 0.209
Table 2: AUROC and PRR scores of UE methods on WebQA and GSM8K. LARS (OOD) denotes that the LARS model is trained with TriviaQA and NaturalQA.
Further, we add 6 non-probability-based UE approaches to our baseline set. Lexical Similarity Fomicheva et al. (2020), is the average of the Rouge-L scores between unique sampled generation pairs to a given question. $\mathbf{p}$(True) Kadavath et al. (2022), a self-check method, asks the model itself if the most likely answer is correct by providing the question, sampled generations, and the answer. SAPLMA Azaria and Mitchell (2023) is a probing-based method that trains the model’s internal representations to predict the correctness of its generation. Eccentricity and Degree Matrix Lin et al. (2024) assesses output consistency using different linear algebraic techniques. Lastly, # Semantic Groups Kuhn et al. (2023) is the number of semantic clusters, as in SE. In all of our experiments, number of sampled generations is 5.
@@
To train the model of the proposed method LARS, we employ train splits of TriviaQA, NaturalQA, and GSM8K. We sample six generations per question, ensuring the most likely generation is included, for each aforementioned model. From these generations, we curate unique QA pairs for calibration data and use GPT-3.5-turbo to evaluate their correctness. Typically, we train distinct LARS models for each model-dataset combination. In some experiments, we merge TriviaQA and NaturalQA per model and train accordingly, which we specify when used. Further details are presented in Appendix D.
@@ 5.2 Main Results
We present the results of our method alongside other baselines in Table 1 and extended results in Appendix C. Notably, LARS significantly enhances the performance of all existing scoring functions across each probability-based UE method, with improvements reaching up to 0.231 AUROC and 0.46 PRR points over LNS. Additionally, LARS boosts the confidence metric to levels comparable with SE. This is particularly important considering the inference costs. Entropy-based methods require multiple output samples (5 in our experiments), which can be computationally expensive in the context of LLMs. Further, SE requires $O(N^{2})$ model passes for semantic clustering, where $N$ is the number of sampled outputs. In contrast, LARS operates with a single pass using a RoBERTa-based model with 125M parameters—a computation level that is negligible compared to models with capacities of 7B parameters or more. Notably, LARS outperforms SAPLMA, which also uses the same amount of supervised data. Additionally, LARS consistently surpasses response clustering methods that require multiple output samples, such as Lexical Similarity, the Number of Semantic Groups, Eccentricity, Degree Matrix, and p(True) method.
@@ 5.3 Out-of-Distribution (OOD) Experiments
First, we investigate how the performance of LARS is affected when the model encounters questions which have a distribution deviating from that of the calibration set. To this end, we conduct tests using WebQA and GSM8K, with LARS models trained on combined TriviaQA and NaturalQA for each distinct chat model. The results are presented in Table 2, and additional results on out-of-distribution (OOD) data generalization are available in Appendix C.3. Impressively, LARS, despite being trained on different datasets, outperforms all other scoring functions across all probability-based UE methods in WebQA, achieving an average improvement of approximately $0.04$ AUROC points. However, in the GSM8K dataset, where the model was trained on a different task, performance degradation becomes significant, highlighting the importance of training LARS on task-specific data for optimal results. This performance gap may be attributed to differences in the nature of the datasets: while TriviaQA answers are primarily composed of entities such as person and organization names, GSM8K primarily involves numerical answers. As a result, calibrating LARS for entity-based answers in TriviaQA makes it less effective for GSM8K, compared to direct calibration on GSM8K itself. Nevertheless, LARS still outperforms other scoring functions in all models except LLama-3-8b, even when not specifically calibrated for the correct dataset.
@@
Next, we analyze how LARS performs when the responses in the calibration set are derived from a different chat model than the one used at test time. Due to space limitations, we provide a subset of the results in Table 3; however, comprehensive results are presented in Appendix C.10. Optimal LARS performance is achieved when the same chat model is used for both training and testing. Nevertheless, OOD model scores still surpass those of baseline scoring functions (see Tables 1 and 6 for baselines), confirming the effectiveness of LARS.
@@
Confidence Llama2-7b 0.858 0.836 0.831
Llama3-8b 0.852 0.874 0.850
Mistral-7b 0.835 0.833 0.852
Entropy Llama2-7b 0.847 0.830 0.827
Llama3-8b 0.852 0.873 0.850
Mistral-7b 0.841 0.841 0.854
SE Llama2-7b 0.850 0.836 0.840
Llama3-8b 0.863 0.872 0.862
Mistral-7b 0.850 0.849 0.859
SentSAR Llama2-7b 0.857 0.841 0.841
Llama3-8b 0.866 0.884 0.863
Mistral-7b 0.851 0.847 0.860
@@
5.4 LARS on Different Languages
To evaluate the performance of LARS and other scoring functions across different languages, we translated the TriviaQA test and calibration datasets into Turkish, German, and Spanish. As shown in Table 4, LARS demonstrates adaptability across languages and outperforms existing scoring functions, showing the importance of calibrating scoring functions for multilingual applications.
Scoring Func. English Turkish German Spanish
LNS 0.747 0.692 0.710 0.701
MARS 0.801 0.695 0.728 0.723
TokenSAR 0.793 0.720 0.758 0.750
LARS 0.864 0.814 0.827 0.835